# Venkata Ramana Makkapati Honda Aircraft Company

### v1

@@ 6 Conclusion
This paper analyzes the “reachable set” of an agent (termed the dependent agent) in instances where it follows another agent (termed the independent agent) using the constant bearing feedback strategy. It is assumed that the speed of the dependent agent is greater than the speed of the independent agent. Theoretical results are presented that analytically characterize the dependent reachable set (DRS) for the instances where the dependent agent’s reachable set does not fully engulf the diameter of the independent agent’s reachable set. It is shown that the theoretical results can be extended to the limiting case where the speeds of both agents are equal. In instances where the dependent agent’s reachable set entirely engulfs either the diameter of the independent agent’s reachable set or the independent agent’s reachable set itself, a tight bound for the DRS is provided along with a hypothesis supported by simulation-based empirical evidence. The study of the DRS for the constant bearing strategy led us to examine a novel optimization problem, which resulted in more empirical evidence-based hypotheses related to the geometry of the constant bearing strategy and the ellipse.

### v2

@@ 6 Conclusion
This paper analyzes the “reachable set” of an agent (termed the dependent agent) in instances where it follows another agent (termed the independent agent) using the constant bearing pursuit strategy. It is assumed that the speed of the dependent agent is greater than the speed of the independent agent. Theoretical results are presented that analytically characterize the dependent reachable set for the instances where the dependent agent’s reachable set does not fully engulf the diameter of the independent agent’s reachable set. It is shown that the theoretical results can be extended to the limiting case where the speeds of both agents are equal. In instances where the dependent agent’s reachable set entirely engulfs the diameter of the independent agent’s reachable set, a tight bound for the dependent reachable set is provided along with a hypothesis supported by simulation-based empirical evidence. The study of the dependent reachable set for the constant bearing pursuit strategy led us to examine a novel optimization problem, which resulted in more empirical evidence-based hypotheses related to the geometry of the constant bearing pursuit strategy and the ellipse.